OpenStudy (anonymous):

Differentiate: y= x^(1-x)

OpenStudy (anonymous):

Use logarithmic properties: \[\Large y(x)=x^{1-x} \\ \Large \ln(y)=(1-x)\cdot \ln x = \ln x - x\ln x \] and then differentiate.

OpenStudy (anonymous):

Is this the answer? x^(1-x) times (-1/x)

OpenStudy (anonymous):